Brightly electrical Company has three factories (𝐴,𝐵, 𝐶) that manufacture desk lamps. A quality control management knows that 35% of lamps are produced by factory 𝐴, 35% by factory 𝐵 and the other 30% by factory 𝐶. He also knows that, generally, 1.5% of desk lamps from factory 𝐴 are defective while 1% 𝑓𝑟𝑜𝑚 𝐵 are defective and 2% from factory 𝐶 are defective.

i.What is the probability that a lamp chosen at random is defective?

ii. If a randomly selected lamp is defective, what is the probability that the lamp was manufactured in factory 𝐶? 


1
Expert's answer
2021-09-27T07:37:03-0400

Let "A" denotes the event that lamp is produced by factory "\ud835\udc34,"

"B" denotes the event that lamp is produced by factory "B,"

"C" denotes the event that lamp is produced by factory "C."

Let "D" denotes the event that lamp is defective.

Given "P(A)=0.35, P(B)=0.35, P(C)=0.3,"

"P(D|A)=0.015, P(D|B)=0.01, P(D|C)=0.02."


i. By the Law of Total Probability


"P(D)=P(A)P(D|A)+P(B)P(D|B)"

"+P(C)P(D|C)"

"=0.35(0.015)+0.35(0.01)+0.3(0.02)"

"=0.01475"

ii. By the Bayes' Theorem


"P(C|D)"

"=\dfrac{P(C)P(D|C)}{P(A)P(D|A)+P(B)P(D|B)+P(C)P(D|C)}"

"=\dfrac{0.3(0.02)}{0.35(0.015)+0.35(0.01)+0.3(0.02)}"

"=\dfrac{0.006}{0.01475}=\dfrac{24}{59}\approx0.40678"